### v3

@@ 7.3 Scorecard mapping
We now make the non-linear tolerance usable in everyday QA. The table below shows how to compute the scorecard fields from the evaluation word count (EWC), the fitted coefficients $\hat{a},\hat{b}$, and the severity-weighted penalty total (APT). The MQM error table itself is unchanged; the only change is that the allowed penalty $E_{\text{allowed}}(x)$ depends on the sample size.
@@
Model coefficients $\hat{a},\hat{b}$ calibrated from tolerance points (see §5.3)
Allowed penalty at size $x$, $E_{\mathrm{allowed}}$ = a*LN(1 + b*EWC)
Quality fraction QFQF $1-\frac{\mathrm{APT}}{E_{\mathrm{allowed}}}$
@@
This keeps the MQM table (error types, error weights, severities and related multipliers $\rightarrow$ APT) unchanged; only the tolerance is dynamic via Equation  (2) and the calibrated mapping stays the same as the linear model’s PT/DPI (Passing Threshold/Defined Passing Interval) presentation.
Figure 7: Example of a real-world non-linear calibrated scoring model for documentation content type (solid yellow curve) plotted against the linear model with 10 allowed penalty points per 1000 words (dashed line). The logarithmic curve is calibrated from two anchor points (not shown in this plot), and therefore represents the actual user tolerance function; the linear line is drawn for comparison only.
Same-curvature retargeting (optional). If you want a new anchor $E(x_{\text{ref}})=E_{0}$ while maintaining curvature, keep $b$ and solve
@@
This re-anchors the tolerance without changing the curve shape.
Micro-range caution. Below $\sim$250 words, deterministic tolerance curves are statistically unreliable; in this sample size range methods of Statistical Quality Control must be used rather than score curves. Quantitatively, if the underlying minor‑error rate is $p$ per word and we observe only $n$ words, the expected count is $np$ with standard deviation $\sqrt{np(1-p)}\approx\sqrt{np}$, so the relative uncertainty scales as $1/\sqrt{np}$ and becomes huge when $np\approx 1$ (e.g., 1 expected error in 200 words at 5/1000).
Worked example (documentation content). Calibrated tolerance: $E(x)=3.688\,\ln(1+0.00288\,x)$. For $x=3000$ words, $E_{\mathrm{allowed}}(x)\approx 8.357$ points. If $\mathrm{APT}=7$, then $QF=1-7/8.357\approx 0.162$. With $\mathrm{PT}=80$, $\mathrm{MSV}=100$ ($\mathrm{DPI}=20$), $\mathrm{OS}\approx 83.25$ and $DM\approx 1.36$ (Pass). If $\mathrm{APT}=9$, then $QF\approx-0.077$, $\mathrm{OS}\approx 78.46$ and $DM\approx-0.64$ (Fail).
@@ 7.4 Implications for CAT Tools and LQA Automation
The change in scoring logic has clear consequences for software systems that implement translation quality evaluation, particularly CAT tools with integrated LQA modules or standalone dedicated LQA platforms.
In the traditional linear model, these tools calculate scores based on a fixed error density relative to a single reference word count. However, with a logarithmic tolerance function, the score calculation must dynamically adjust based on the actual sample size. This means LQA modules and dashboards need to:
  • Replace static pass/fail thresholds based on a single anchor point with dynamic tolerance functions, $E(x)$, that adjust based on the actual word count ($x$).
  • Visualize quality scores against a curved baseline that represents the tolerance function, not a simple flat line.
  • Provide evaluators with interactive feedback showing how close a sample’s error count is to its unique, size-adjusted pass/fail threshold.
This functionality is particularly valuable for machine translation, post-editing, and AI-generated content, as evaluation samples from these sources often vary widely in word count.
7.5 Improvement of a Unified Calibrated Quality Scale
One major challenge in TQE is ensuring that quality scores remain human-readable, interpretable, and comparable across different types of content, language pairs, and evaluation sample lengths. The linear calibrated model is a significant improvement over using raw scores with very small passing intervals, but even the linear calibrated model can distort comparability when applied to large or small samples.
With a non-linear calibrated scoring model, it becomes feasible to normalize all evaluations to a unified, user-specific quality scale. Since the logarithmic tolerance accounts for sample-size effects, and calibration maps results to a common 0–100 (PT/MSV) scale, all evaluations land on the same company-wide quality scale—so scores from jobs of different lengths, language pairs and content types are directly comparable. For example, replicating the English–German marketing case from Table 3 in [GladkoffEtAl2024], a 5,000-word translation with 23 minor errors passed the linear model, but was rejected by human reviewers. The logarithmic model set a threshold of 16 errors—correctly matching the expert decision.
This normalization also improves the validity of quality measurement results analytics across a portfolio of projects, enabling better tracking of supplier performance, trend analysis, and root-cause investigations.
@@ 7.6 Applications Beyond Human Translation
While the immediate application of this model is human translation quality evaluation, the underlying principles are equally relevant to assessing AI-generated content. Large language model (LLM) outputs—such as summaries, answers, or document translations—are increasingly being evaluated using scoring methods similar to those applied for human translation quality evaluation [han2024neural].
The model based on non-linear error tolerance provides a fairer and more realistic baseline for evaluating content. It avoids the common pitfalls of unduly penalizing shorter segments with one mistake or overlooking the accumulated minor issues in long, highly fluent AI-generated texts.
As organizations adopt more automated content generation, the ability to evaluate quality in a human-aligned, perceptually realistic manner becomes essential. The proposed scoring model provides a foundation for doing so.
@@ 7.7 Document-Level Evaluation and Future Automation
Translation quality never exists in a vacuum: every segment lives within the cohesive fabric of a document. A translation sample is therefore not a roll of toilet paper to be inspected centimeter by centimeter; its quality must be judged in the context of the larger text. The larger the sample relative to the document, the more faithfully that judgment reflects holistic properties such as cohesion, coherence, register, and terminological consistency.
Small extracts can still provide a rapid signal, but offer only a limited view of document quality. They cannot capture discourse-level phenomena, and—because error density interacts non-linearly with length—linear metrics break down outside their narrow reference window.
The MQM-based Non-Linear Calibrated model solves this by remaining simple and linear in the meso-range yet extending smoothly into the macro-range. As long as the evaluation sample is sufficiently large relative to the document, as recommended by the Multi-Range Framework [GladkoffEtAl2024]—it yields scores that are statistically reliable and consistent with the judgment of experts at the document level.
Of course, analytic TQE becomes increasingly expensive at large evaluation sample sizes. No current AI system can yet perform a complete analytic evaluation, but future models may acquire the ability to conduct formal TQE (full analytic MQM evaluation) over long contexts (document-scale samples), reliably capturing local errors (segment-level) and holistic properties (cohesion, coherence, terminology/style consistency, discourse). The logarithmic model proposed here is expressly designed to serve as the computational backbone for such future automation, ensuring that automated scoring remains perceptually valid across the full spectrum of document lengths.
@@ 7.8 Summary
In practical terms, the transition to a non-linear scoring model means:
  • Greater alignment between future automated document-level scoring systems and expert human judgment,
  • Improved fairness and consistency across evaluation samples of different lengths,
  • More intelligent quality evaluation systems producing evaluations closer to human perception,
  • Stronger foundations for evaluating both human and AI-generated translation.
This model equips both translation professionals and tool developers with a framework that respects the cognitive and perceptual realities of language evaluation, enhancing decision-making at every level of the translation quality pipeline. The proposed non-linear model captures how translation quality is actually perceived by humans-especially when evaluation sample sizes vary. It is simple to implement, easy to calibrate, and theoretically grounded. By adopting non-linear scoring, evaluation systems can better align numeric scores with human judgment, increasing the reliability and fairness of translation quality evaluations, as well as increasing the reliability and fairness of quality evaluations of translations created by humans or AI-generated texts.
@@ 8 Limitations and Future Work
The following limitations and research directions qualify the benefits summarized in Section 7.8 and the conclusion. While the findings and proposed non-linear quality evaluation model presented in this paper are supported by both empirical observation and theoretical grounding, there are important limitations that must be acknowledged. Addressing these will be essential for refining the model and validating its broader applicability.
8.1 Limited Survey Group Size
We conducted structured interviews with quality managers at three very large enterprise institutes, and the results are corroborated by their internal practices. Despite the scale and maturity of these organizations’ multilingual translation programs, the interview sample is still small. Additional evidence was drawn from our own practice, where we exclusively use a non-linear calibrated scoring model due to its universality and accuracy. However, the tolerance data we used was gathered through expert elicitation and professional experience, rather than under blind experimental conditions—introducing subjectivity and potential sources of bias, including response bias.
Further research should involve a broader range of organizations—including LSPs, in-house localization teams, and public-sector translation departments—across different domains and language pairs. Ideally, future studies should also include controlled experiments where evaluators independently assess translation quality using both holistic judgment and formal scoring models, enabling direct comparison between model predictions and actual pass/fail outcomes.
8.2 Content-Type Specific Effects
Different content types such as marketing, legal, technical, and UI strings, may have different quality expectations and error tolerance. Tolerance growth is assumed to follow the same logarithmic form $E(x)=a\,\ln(1+bx)$; the two coefficientsaa (overall scale) and $b$ (curvature)—may differ by user or use case.
We found that a single logarithmic formula performs well across a wide range of content types, suggesting that the logarithmic rule is quite universal—even with fixed parameters. However, this observation requires further validation through broader field data. A valuable extension of this work would be to investigate whether different genres (content types) exhibit distinct curve shapes—for example, steeper or flatter logarithmic profiles—or whether certain domains deviate significantly from the general trend. Such an analysis would help determine whether the proposed model serves as a general framework, as our experience suggests, or represents just one of several viable scoring paradigms. We hypothesize a soft saturation at very large evaluation wordcount due to document‑level coherence expectations; future work can test whether $E(x)$ bends further (e.g., log‑log or logistic) at book/manual length.
8.3 Severity Weighting and Mixed Error Types
This paper concentrates on modeling tolerance in terms of penalty points, each of which typically represents one minor error. In practice, quality evaluations combine several error severities (minor, major, critical) penalized by specific multipliers and sometimes further scaled by different error-type weights applied for different content types.
Severity scaling, i.e., fixed severity weights such as minor/major/critical multipliers, is incorporated by assigning a major or critical error a fixed multiple of the minor penalty points. This is separate from any error-type weights (e.g., terminology vs. grammar) a metric may also define. While a logarithmic curve fits minor-error tolerance well—minor errors being the most frequent and thus the most statistically stable—it may be less suitable for major errors, which occur far less often. For these rarer severities, an integer-valued sequence may be preferable, as indicated by Institution 3’s data in Fig. 3. That sequence follows a loosely logarithmic trend but would require explicit rounding rules.
Future work should examine the form of such integer sequences, quantify how far they depart from a pure logarithmic curve, and determine the parameters that best describe severity scaling across different content types and languages.
8.4 Inter-Rater Reliability, IRR
When a linear model is used, the variability among human raters can be partly attributed to the fact that the model is only accurate at its anchoring point and loses precision elsewhere. In contrast, a calibrated non-linear model better reflects human judgment across a wide range of evaluation sample sizes. As a result, Inter-Rater Reliability (IRR) should increase significantly when using a non-linear calibrated scoring model, as it provides consistent and comparable measurements across different evaluation sample sizes, languages, and content types. More research is needed to quantify the improvement in inter-annotator agreement and to determine how much such a model reduces variance compared to traditional linear approaches in real-world annotation settings.
@@ 8.5 Extension to AI Evaluation Benchmarks
This paper also hypothesizes that the same principles of non-linear error perception apply to AI-generated text evaluation, including summarization, dialogue, and document translation by large language models. However, this claim has yet to be formally validated. AI benchmarks such as WMT [^10] https://machinetranslate.org/wmt shared task series [kocmi-etal-2025-findings] which rely on reference-based scoring or raw error density metrics can be improved by introducing calibrated non-linear models which could improve their alignment with human preferences.
Future work should involve testing the non-linear model as a normalization layer in LLM evaluation tasks, particularly those involving paragraph- or document-level generation, where holistic perception plays a major role.
A further implication is that a calibrated non-linear scoring model could eventually serve as a reference framework for future AI-based document-level evaluations, which currently provide holistic but qualitative judgments without a quantitative calibration.
@@ 9 Conclusion
Translation quality evaluation plays a critical role in ensuring that multilingual content meets the linguistic, technical, and functional standards required for global communication. While current industry practices predominantly rely on linear scoring models—extrapolating error thresholds proportionally from fixed reference wordcounts—this paper builds on empirical and practitioner evidence showing that linear models do not consistently align with human evaluation, motivating the adoption of non-linear alternatives.
This paper presents empirical data from experienced quality managers at large institutional translation service users, all of whom consistently demonstrated that their internal sense of acceptable error count rises non-linearly with sample size. This effect was evident across organizations, content types, and error categories and severities. This paper further supports these observations with theoretical underpinnings from psychophysics (via the Weber–Fechner law) and cognitive science (via Cognitive Load Theory), both of which explain why perceived quality and impact of errors do not scale linearly with text length.
To address this disconnect between the existing scoring systems and human judgment, this paper proposes a calibrated logarithmic scoring model that reflects the genuine growth of error tolerance. This model is mathematically simple, empirically grounded, and easily adaptable to specific user requirements. It provides a flexible, yet rigorous alternative to traditional linear error density calculations. Because it retains a simple two-parameter form—and requires only a one-dimensional solve for calibration—the model is straightforward to integrate into CAT tools, dashboards, and analytics, while providing a unified way to reflect how perceived impact scales with error accumulation.
The implications of adopting this model are wide-ranging. Quality managers can gain a more accurate tool for defining passing thresholds in jobs of variable lengths. CAT tools and LQA systems and tools can offer context-aware scoring and fairer quality metrics. Evaluators benefit from consistency that reflects actual perception, and the same principles can be extended to evaluating AI-generated language output. While further research is needed to validate and refine this model, particularly across languages, domains, content types, the fundamental insight seems to be clear: human perception of translation quality is inherently non-linear. By embracing this fact, the translation industry can take a critical step toward a more accurate, fair, and human-perception-aligned quality evaluation.
Finally, this paper reiterates the warning from [GladkoffEtAl2024]: below the sample size of approximately 250 words, no deterministic tolerance curve is statistically reliable. In such micro-range cases, binomial risk-based Statistical Quality Control (SQC) remains the only viable method. [^11] Even with SQC, micro-range evaluations cannot provide statistically reliable estimates of the underlying population quality, because the confidence intervals around error density are too wide to be meaningful at such small evaluation sample sizes. SQC explicitly departs from the idea of estimating population quality and instead evaluates the producer’s risk (the risk of wrongly rejecting acceptable output) and the consumer’s risk (the risk of wrongly accepting defective output). In this way, SQC reframes quality control for very small samples: It does not reduce statistical uncertainty, but it makes risks explicit and quantifiable for decision-making. This study consolidates the multi-range theory findings by demonstrating that the same non-linear curve can be independently derived from psychophysical theory and empirical data, thus closing the theoretical gap identified in that earlier work.

### v4

@@ 7.3 Scorecard mapping
Table 2 summarizes the computations required to populate an operational scorecard using the evaluation word count (EWC), fitted coefficients $\hat{a},\hat{b}$, and severity-weighted penalty total (APT). Importantly, the MQM annotation process and error-type/severity schema remain unchanged; only the tolerance threshold becomes length-dependent via Equation (2).
@@
Model coefficients $\hat{a},\hat{b}$ calibrated from tolerance points (Section 5.3)
Allowed penalty $E_{\mathrm{allowed}}(x)$ = a*LN(1 + b*EWC)
Quality fraction $QF(x)$ $1-\frac{\mathrm{APT}}{E_{\mathrm{allowed}}}$
@@
Figure 7 illustrates a real-world non-linear tolerance curve calibrated from institutional anchor points, alongside the corresponding anchored linear rule for comparison.
Figure 7: Example non-linear tolerance curve (solid) compared with an anchored linear rule (dashed). The logarithmic curve is calibrated from tolerance points and represents institutional tolerance as a function of sample size; the linear curve is shown for reference.
Same-curvature re-anchoring (optional). To impose a new anchor value $E(x_{\text{ref}})=E_{0}$ while preserving curvature, fix $b$ and solve
@@
Micro-range caution. For sample sizes below approximately 250 words, deterministic tolerance curves are statistically unstable, and Statistical Quality Control (SQC) procedures are recommended instead. Intuitively, if the underlying minor-error rate is $p$ per word and only $n$ words are observed, the expected count is $np$ with standard deviation $\sqrt{np(1-p)}\approx\sqrt{np}$, so relative uncertainty scales as $1/\sqrt{np}$ and becomes large when $np\approx 1$.
Worked example (documentation content). Calibrated tolerance: $E(x)=3.688\,\ln(1+0.00288\,x)$. For $x=3000$ words, $E_{\mathrm{allowed}}(x)\approx 8.357$ points. If $\mathrm{APT}=7$, then $QF\approx 0.162$. With $\mathrm{PT}=80$, $\mathrm{MSV}=100$ ($\mathrm{DPI}=20$), $\mathrm{OS}\approx 83.25$ and $DM\approx 1.36$ (Pass). If $\mathrm{APT}=9$, then $QF\approx-0.077$, $\mathrm{OS}\approx 78.46$ and $DM\approx-0.64$ (Fail).
@@ 7.4 Implications for CAT tools and LQA automation
Adopting a logarithmic tolerance function has direct implications for software systems implementing LQA workflows, including CAT tools with integrated quality modules and standalone QA platforms. Under proportional scaling, thresholds are defined using a constant error density relative to a fixed reference length; under the proposed model, tolerance must be computed dynamically as a function of the observed evaluation word count.
In practice, this requires tooling to:
  • replace static pass/fail thresholds with a length-dependent tolerance function $E(x)$;
  • visualize quality relative to a curved tolerance baseline rather than a proportional reference line; and
  • report decision margins (e.g., $DM=E_{\mathrm{allowed}}-\mathrm{APT}$) to communicate proximity to the pass/fail boundary.
These adaptations are particularly relevant in workflows involving heterogeneous sample sizes, such as machine translation, post-editing, and AI-assisted content generation.
7.5 Toward a unified calibrated quality scale
A persistent challenge in analytic TQE is maintaining interpretability and comparability across content types, language pairs, and evaluation sample lengths. Calibrated scoring improves over raw penalty rates by mapping results to a stable reporting scale; however, proportional length scaling can still distort comparability when applied outside its effective range.
By incorporating length-dependent tolerance via $E(x)$, non-linear calibrated scoring supports a unified, institution-specific 0–100 scale for heterogeneous evaluation sizes. This improves comparability across jobs and enables more reliable aggregation for analytics such as trend monitoring, supplier benchmarking, and root-cause analysis.
As a concrete example, [GladkoffEtAl2024] reports a case in which a 5,000-word English–German marketing translation containing 23 minor errors passed under linear scaling but was rejected by expert reviewers. The logarithmic model yields a substantially lower tolerance threshold in this regime (16 minor-error equivalents), consistent with the human accept/reject decision.
@@ 7.6 Applications beyond human translation
Although the model is motivated by human translation evaluation practices, the same length-dependent tolerance considerations apply to AI-generated text. Large language model (LLM) outputs (e.g., summaries, answers, or translations) are increasingly evaluated using procedures adapted from TQE and LQA [han2024neural]. Non-linear tolerance modeling provides a more stable baseline for such evaluation by reducing over-penalization of short segments and preventing under-penalization of accumulated minor issues in long outputs.
@@ 7.7 Document-level evaluation and future automation
Document-level quality assessment requires integrating local defects with discourse-level properties such as cohesion, register consistency, and terminology management. Small samples can provide rapid quality signals, but they provide only partial visibility into document-level phenomena and may lead to unstable conclusions when length effects are ignored.
The MQM-based non-linear calibrated model addresses this by maintaining linear behavior near its reference regime while extending smoothly to longer samples via a single tolerance function. Provided the evaluation sample is sufficiently large to yield statistically stable estimates (as recommended by the Multi-Range framework [GladkoffEtAl2024]), the resulting scores remain consistent with document-level expert judgments.
Finally, although current automated systems do not yet perform complete MQM-style analytic evaluation over long contexts, future evaluation models may increasingly combine local error detection with document-level assessment. In such settings, a length-dependent tolerance function provides a principled backbone for ensuring that automated scoring remains comparable and perceptually aligned across a wide range of evaluation lengths.
@@ 7.8 Summary
In practical terms, transitioning to a non-linear tolerance model enables:
  • improved consistency of acceptability decisions across variable-length evaluation samples;
  • more interpretable and equitable calibrated scores on a unified reporting scale;
  • simpler integration into CAT and LQA tooling via dynamic tolerance computation; and
  • better alignment with document-level evaluation requirements for both human and AI-generated content.
Overall, the proposed approach operationalizes empirically supported non-linear tolerance scaling in a form that is simple to calibrate, straightforward to implement, and compatible with existing MQM-based workflows.
@@ 8 Limitations and Future Work
This section qualifies the practical benefits summarized in Section 7.8 and identifies directions for further validation and extension. Although the proposed non-linear tolerance model is supported by both expert elicitation and theoretical motivation, several limitations constrain the scope of the current evidence base.
8.1 Sample size and elicitation design
The empirical evidence in Section 3 is based on structured elicitation with Quality Managers from three large institutional translation programs. While these organizations operate mature, high-volume multilingual workflows, the number of participating institutions (and respondents) remains limited. Moreover, tolerance judgments were elicited rather than measured in blinded controlled experiments, which may introduce subjectivity and potential response bias.
Future work should expand the sample to include a broader range of stakeholders—e.g., LSPs, in-house localization teams, and public-sector translation units—across domains and language pairs. Controlled studies could further strengthen validity by comparing (i) holistic accept/reject decisions and (ii) analytic MQM-based scoring outcomes under alternative tolerance functions, enabling direct assessment of predictive alignment and decision consistency.
8.2 Content-type and domain effects
Quality expectations vary by content type (e.g., marketing, legal, technical documentation, UI strings). The proposed model assumes a common functional form $E(x)=a\ln(1+bx)$, with content- or user-specific variation captured by the parameters $a$ (scale) and $b$ (curvature). Our operational observations suggest that a single logarithmic form can perform robustly across heterogeneous content types, but broader field evidence is required.
An important extension is to test whether particular domains exhibit systematically different curvature profiles (e.g., steeper for high-risk instructional content or flatter for low-stakes internal communication), and whether additional saturation occurs for very large evaluation sizes (e.g., book- or manual-length contexts). If such effects are observed, alternative sublinear families (e.g., log–log, logistic saturation, or piecewise concave forms) may be warranted.
8.3 Severity weighting and mixed error distributions
This paper models tolerance in severity-weighted penalty points, where one point is typically interpreted as one minor-error equivalent. In operational settings, evaluations include mixed severities (minor/major/critical) with fixed multipliers and may also include error-type weights. Although the logarithmic tolerance curve captures minor-error tolerance well—minor errors being the most frequent and thus statistically most stable—rarer severities may require discrete (integer) decision rules or severity-specific tolerance sequences.
Future work should evaluate how tolerance behaves for major and critical errors under variable-length sampling. This includes establishing principled rounding and decision rules, testing whether severity-specific tolerance sequences remain approximately sublinear, and assessing whether severity multipliers interact with length-dependent tolerance in ways that call for multi-parameter extensions.
8.4 Inter-rater reliability
A key motivation for length-aware tolerance is to reduce systematic disagreement caused by applying proportional extrapolation outside its effective reference region. The proposed model is expected to improve inter-rater reliability (IRR) by providing more consistent pass/fail thresholds across varying sample sizes. However, the present work does not yet quantify IRR improvements under controlled rater studies.
Future work should measure IRR (e.g., using Krippendorff’s $\alpha$, Cohen’s $\kappa$, or intra-class correlation, depending on task design) under both linear and non-linear scoring regimes. Such studies could test whether a length-dependent tolerance model reduces variance in borderline cases and improves agreement across evaluators with different experience levels.
@@ 8.5 Extension to AI evaluation benchmarks
We hypothesize that similar length-dependent tolerance effects may apply to the evaluation of AI-generated text (including document translation, summarization, and dialogue). This claim remains to be validated empirically. Benchmarking settings—including shared-task evaluations in MT such as WMT [kocmi-etal-2025-findings]—often rely on reference-based metrics or length-normalized error densities; incorporating calibrated non-linear normalization may improve comparability across variable-length outputs and strengthen alignment with human preferences in document-level evaluation.
Future work should test the proposed model as a normalization layer for paragraph- and document-level evaluation in both MT and broader LLM generation tasks, including analysis of robustness across domains and prompt conditions. A longer-term direction is to investigate whether calibrated tolerance functions can serve as quantitative scaffolding for AI-assisted document-level quality assessment systems that currently produce holistic but weakly calibrated judgments.
@@ 9 Conclusion
This paper addressed the length-dependence of acceptability thresholds in MQM-based analytic Translation Quality Evaluation (TQE). While proportional (linear) error scaling is widely used in practice, our results indicate that it can produce systematic bias when evaluation sample sizes deviate from the reference word count.
First, we provided empirical evidence from three large institutional translation programs showing that tolerable error counts increase with sample length sublinearly, yielding concave tolerance profiles consistent with logarithmic growth. Second, we linked this behavior to established theoretical accounts from psychophysics (compressive perceptual scaling) and Cognitive Load Theory (accumulated processing disruption), which together motivate non-linear integration of errors over increasing text length. Third, we proposed a calibrated two-parameter tolerance function $E(x)=a\ln(1+bx)$ and showed how it can be anchored from tolerance points using a simple one-dimensional calibration procedure. Fourth, we demonstrated that the model can be integrated into existing MQM pipelines with minimal change: the MQM annotation process and penalty computation remain unchanged, while a constant length-normalized tolerance is replaced by a length-dependent threshold function.
In addition, we derived a near-reference fidelity interval that characterizes when a linear approximation remains sufficiently accurate (e.g., within $\pm 20\%$), providing a practical guideline for model selection in operational deployments. Finally, we reiterate that for very short samples (approximately $\lesssim 250$ words), deterministic tolerance curves are statistically unstable, and risk-based Statistical Quality Control (SQC) remains the appropriate framework for decision-making under high sampling uncertainty [GladkoffEtAl2024].